# MatSynth: A Modern PBR Materials Dataset

### v1

@@ Abstract
We introduce MatSynth, a dataset of $4,000+$ CC0 ultra-high resolution PBR materials. Materials are crucial components of virtual relightable assets, defining the interaction of light at the surface of geometries. Given their importance, significant research effort was dedicated to their representation, creation and acquisition. However, in the past 6 years, most research in material acquisiton or generation relied either on the same unique dataset, or on company-owned huge library of procedural materials. With this dataset we propose a significantly larger, more diverse, and higher resolution set of materials than previously publicly available. We carefully discuss the data collection process and demonstrate the benefits of this dataset on material acquisition and generation applications. The complete data further contains metadata with each material’s origin, license, category, tags, creation method and, when available, descriptions and physical size, as well as 3M+ renderings of the augmented materials, in 1K, under various environment lightings. The MatSynth dataset is released through the project page at: https://www.gvecchio.com/matsynth.
@@ 1 Introduction
Materials are a key component of the rendering pipeline, describing the appearance of objects, for example in terms of colours, shininess or metallicity. The design of high quality virtual materials is a complex endeavour, usually requiring significant artistic expertise [29]. To reproduce existing objects and surfaces, material acquisition and generation have been active research fields for decades [14]. More recently, the field turned to Machine Learning to tackle the ill-posedness of few images (lightweight) material acquisition [23, 5], and to explore new opportunities in material generation[18, 49]. Since 2018, most of the field relies on a single dataset [5] or, when possible, on company-owned large material libraries [1, 27, 49, 43]. As company owned libraries are typically orders of magnitude bigger than what is publicly available, they enable very high quality results and experiments with higher resolutions. These materials are however not availabe to the community, underlining the need for larger, more diverse, public datasets to enable further research. Such material data is also particularly useful for synthetic data generation, for example for object acquisition [24, 8], material based selection [37], and numerous applications further from material understanding and authoring [39, 33, 41].
To bridge the gap between the private and public datasets, we present MatSynth, a dataset of $4,069$ tileable materials, available in 4K. Alongside the material properties themselves, we render $3,417,960$ renderings with various scales cropping, rotation and environment illuminations to facilitate future learning. The materials all come with permissive licences and were sourced from various libraries [31, 36, 2, 4, 40] and checked for duplication among the newly gathered data and against previous datasets [5]. We further ensure that all materials are represented by the same set of parameters and that they are tileable. In particular each material is represented by 7 textures/maps: Base color, Diffuse Normal, Height, Roughness, Metallic and Specular, ensuring compatibility with both Diffuse/Specular (e.g. [5]) and Base Color/Metallic (e.g. [27]) workflows. Some materials in our dataset also have an opacity map to handle, for example, surfaces with holes. Finally, we process the previously available dataset to enrich it with metallic, base color and height properties, facilitating their combinations for a total of $\sim 6,000$ materials available for research with clear licenses attached.
To validate the benefit of this additional data, we train recent methods with publicly available implementation, and evaluate both qualitatively and quantitatively the results quality for different single image material capture methods [5, 42] and material generation [44], showing that the additional data enables higher quality acquisition and improved diversity in the generation results.
@@ 2.1 Existing dataset
The only large-scale, publicly available, spatially-varying material dataset contains 1628 distinct, pre-augmentation materials [5] with a research-only license. 512x512 [6] and 2K [7] versions of the materials were later published. It is based on random variations from a set of 155 procedural materials gathered from Allegorithmic Substance Share in 2017, distributed in 9 categories: paint(6), plastic(5), leather(13), metal(35), wood(23), fabric(6), stone(25), ceramic tiles(29), ground(13). Other datasets used in the field contain 100 and 62 measured BRDFs [28, 10] 170 SVBRDFs licensed from VRay[23, 45] or more recently 211 materials from online libraries [17]. Adobe 3D Assets [1]’s $11,000+$ procedural assets have been used to unlock significant quality improvement in the methods that could benefit from it for acquisition, generation, retrieval and captioning [49, 27, 43, 9]. This data is however difficult to license for Machine Learning, limiting its use to company internal users. In contrast, we propose a systematic gathering of non-duplicate, tileable, publicly available materials with permissive licenses from various online libraries, assembling $4,000$+ materials and their metadata, significantly increasing the variety and amount of ressources available for material-related research.
@@ 3 The Dataset
The MatSynth dataset is designed to support modern, learning-based techniques for a variety of material-related tasks including, but not limited to, material acquisition, material generation and synthetic data generation e.g. for retrieval or segmentation. It improves on existing available datasets by providing a large collection ($\sim 2.5$ times larger than previously available datasets) of non duplicate, high-quality, high-resolution realistic materials alongside useful metadata. Each material in the dataset is represented by a set of reflectance maps (basecolor, diffuse, normal, height, roughness, metallic and specular), complemented by fine-grained annotations including the source and link to the asset, its license and original author (when available), tags, the creation method, whether is is stationary or not, a timestamp for versioning, a description (when available), and the material physical size (when available).
@@ 3.3 Data Processing
Finally, we automatically compute an estimate of the correct displacement factor for each material, by optimizing a displacement factor parameter to match the normal vectors derived from the scaled height map to the ground truth normal map. This ensures a good match between the normal and the displacement scale when rendering. We provide the correctly scaled height map, as well as the original and the scaling parameters (scale factor and mid value compatible with the Blender ”Displacement” node).
@@ 3.4 Data Augmentation
Rotation and crops and illumination The collected 4069 materials are augmented by rotation and cropping. In particular, we extract 168 unique crops for each material, at different scales. The crops are extracted in the following way. The original material is rotated 8 times ($0\degree$, $45\degree$, $90\degree$, $135\degree$, $180\degree$, $225\degree$, $270\degree$, $315\degree$)–for angles that are not multiple of $90\degree$, we first tile the material, apply the rotation, and finally crop the center of the material at 4K resolution. For each rotation, we extract 16 non-overlapping crops at 1K resolution, 4 at 2K resolution, 1 at 4K resolution, for a total of 21 crops per rotation. All crops are resized to a resolution of $1024\times 1024$ pixels. Each crop is rendered with under different environment illuminations (2 outdoors, 2 indoors, and 1 studio light), for a total of 840 renders per material. For each render, a random rotation is applied to the environment map. The resulting dataset consists of 683,592 material crops and 3,417,960 renders. Fig. 2 shows the five renderings under different environment illumination for 4 different materials from the dataset. We include the 5 environment maps used for rendering the dataset in the supplemental materials.
@@
Rendering We render each material displacing the mesh using the available Height map, thus introducing realistic cast shadows in the rendering. However, to maintain a pixel match with the maps, required for supervised SVBRDF estimation tasks, we opt for an orthographic rendering. This introduces yet another issue, as specular reflections are lost on perfectly flat surfaces due to rays being parallel when rendering with an orthographic camera. To restore specular reflection while keeping pixel match and cast shadows, we split the rendering in two separate passes. Similar to the data used to train ControlMat [43], we first render the diffuse component of the material using an orthographic camera with displacement enabled, we then render the specular component using a perspective camera, while disabling the displacement. The two renders are then added back to get the final render. Fig. 3 shows the output of the two render passes and the final result. For each render we provide the final image, as well as the intermediate passes.
@@
Figure 3: Render samples using the two-pass strategy. This strategy ensures that the maps and the rendering are well aligned, avoiding parrallax effect caused by perspective, but preserving specular highlights.
@@ 3.6 Dataset Statistics
We report here statistics of our dataset in terms of categories, tags, creation methodology, stationarity and number of asset per source.
Categories. We report the category histogram in Fig. 4, showing a wide diversity of appearances, and compensating for some gaps in the distribution of previously available materials. Tags. We have $21,737$ tags from $1,239$ unique tags associated with the materials. For example, the top five most represented tags are floor(924), dirty(543), man (504), wood (474), brown (444) and wall(424). Each material has between 0 (371 assets didn’t have associated tags) and 20 tags with an average of 5.74, a standard variation of 4.35 and a median of 5.
@@
Stationarity. Our dataset contains 3096 stationary and 973 non stationary materials. We compute stationarity by extracting 8 random crop of the basecolor and height map and comparing the CLIP cross-similarity matrix. If more than 12.5% of the cross similarities are below a threshold (we empirically found 0.9 to work well), the material is not considered stationary.
@@
Table 1: Quantitative results. We report the RMSEnormal-$\downarrow$ between predicted and ground-truth maps and renderings (averaged over renderings under 5 environment lightings), except for Normal maps for which we report the cosine errornormal-$\downarrow$. We see that the additional data provided by MatSynth enables higher quality material acquisition both in terms of parameters and renderings comparisons.
@@
Table 2: Perceptual quantitative results. We report here the perceptual metrics for the Renderings, rendered under 5 environment lightings (which can be intepreted as natural images) SSIMnormal-$\uparrow$ and LPIPSnormal-$\downarrow$. Similar to the RMSE evaluation, additional data significantly improves the rendering matching’s quality.
@@ 4.1 Test dataset
We build a static test set, with the aim of simplifying comparison of future work, by handpicking 5 unique samples for each cathegory. We further ensure that none of the chosen samples is a simple variation (e.g.: color or scale) of other items in the training set to avoid any bias. We merge this dataset, consisting of 65 materials, with the test set from Deschaintre et al. [7], for a total of 89 test materials.
@@ 4.2 Evaluation Methodology
Evaluation is carried on on two main materials–related tasks: estimation and generation. For both tasks, we look at methods with publicly available implementations. In particular, we evaluate Deschaintre et al. [5] and SurfaceNet [42] for materials capture performance, and MatFuse [44] for material generation. In both cases, we focus on the diffuse/specular rendering workflow for an easier comparison with the current state of the art. We additionally propose an evaluation of SurfaceNet trained using the basecolor/metallic workflow in Supplemental material, to set a new baseline for future work shifting to this workflow.
@@ 4.3 Training Strategy
All methods are trained on a nVidia RTX 4090 GPU.
4.3.1 Deschaintre et al. (2018)
We train the method using the author’s publicly available Tensorflow implementation using a combination of our dataset and the original one. Following the original paper we train the model for $400,000$ steps (72 hours) with a batch size of 8 and a learning rate of $2^{-5}$.
4.3.2 SurfaceNet
Similarly to the training of Deschaintre et al. [5], we use the author’s original Pytorch implementation and the combined dataset. We also follow the original paper and train the model for $400,000$ steps (96 hours) with a batch size of 8 and a learning rate of $2^{-5}$. We include the adversarial discriminator after $100,000$ steps.
4.3.3 Diffusion Model
We evaluate the generation quality improvement provided by our dataset by fine-tuning the recent Matfuse [44]. We use the same multi-encoders compression model, pre-trained on the Deschaintre et al. [5]’s dataset, but fine-tune the latent diffusion model itself. In particular, the diffusion model is fine-tuned for 150,000 iterations with a batch size of $20$ and a learning rate of $10^{-4}$. We use a linear learning rate warm-up until step 10,000. For inference, following the original paper, the denoising is done using a DDIM sampling schedule [38] with $50$ timesteps.
@@ 4.4 SVBRDF Estimation Results
Figure 7: Comparison on Material Generation. We show renderings of 15 randomly generated materials, both with and without our dataset. These results are directly sampled from the network and have not been filtered. As expected, the additional data helps achieve better and more diverse results.
@@ 4.5 Materials Generation Results
We evaluate material generation using MatFuse [44], trained on Deschaintre et al. [5]’s dataset only against a version fine-tuned with MatSynth. As material maps have different distributions than natural images, we report FID on renderings under environment lighting. For fairness we only use Deschaintre et al. [5]’s dataset to compute the Ground-Truth distribution and compute the FID for each method using $1,000$ random samples. Before fine-tuning we find a FID score (lower is better) of $239.9$ against $210.3$ after fine-tuning.
We further evaluate the benefit of the additional data for diversity and quality. We show in Table 3 the percentage of samples belonging to different classes, and show a qualitative comparison in Fig. 7. Both evaluations show a wider results diversity and Fig. 7 shows that the generation results are more realistic after fine-tuning.
W/o fine-tuning W/ fine-tuning
Wood 36.1% 25.72%
Ceramic 1.1% 12.51%
Stone 21.5% 14.8%
Metal 3.45% 8.4%
Fabric 0.33% 7.34%
Ground 1.8% 16.2%
Others 35.72% 15.03%

### v3

@@ Abstract
We introduce MatSynth, a dataset of $4,000+$ CC0 ultra-high resolution PBR materials. Materials are crucial components of virtual relightable assets, defining the interaction of light at the surface of geometries. Given their importance, significant research effort was dedicated to their representation, creation, and acquisition. However, in the past 6 years, most research in material acquisition or generation relied either on the same unique dataset, or on company-owned huge library of procedural materials. With this dataset, we propose a significantly larger, more diverse, and higher resolution set of materials than previously publicly available. We carefully discuss the data collection process and demonstrate the benefits of this dataset for material acquisition and generation applications. The complete data further contains metadata with each material’s origin, license, category, tags, creation method, and, when available, descriptions and physical size, as well as 3M+ renderings of the augmented materials, in 1K, under various environment lightings. The MatSynth dataset is released through the project page at: https://www.gvecchio.com/matsynth.
@@ 1 Introduction
Materials are a key component of the rendering pipeline, describing the appearance of objects, for example in terms of colors, shininess or metallicity. The design of high-quality virtual materials is a complex endeavor, usually requiring significant artistic expertise [30]. To reproduce existing objects and surfaces, material acquisition and generation have been active research fields for decades [14]. More recently, the field turned to Machine Learning to tackle the ill-posedness of few images (lightweight) material acquisition [23, 5], and to explore new opportunities in material generation[18, 50]. Since 2018, most of the field has relied on a single dataset [5] or, when possible, on company-owned large material libraries [1, 28, 50, 44]. As company owned libraries are typically orders of magnitude bigger than what is publicly available, they enable very high quality results and experiments with higher resolutions. These materials are, however, not available to the community, underlining the need for larger, more diverse, public datasets to enable further research. Such material data are also particularly useful for synthetic data generation, for example, for object acquisition [24, 8], material-based selection [38], and numerous applications further from material understanding and authoring [40, 34, 42].
To bridge the gap between private and public datasets, we present MatSynth, a dataset of $4,069$ tileable materials, available in 4K. Alongside the material properties themselves, we render $3,417,960$ renderings with various scales cropping, rotation, and environment illuminations to facilitate future learning. The materials all come with permissive licences and were sourced from various libraries [32, 37, 2, 4, 41] and checked for duplication among newly gathered data and against previous datasets [5]. We further ensure that all materials are represented by the same set of parameters and that they are tileable. In particular, each material is represented by 7 textures/maps: Base color, Diffuse, Normal, Height, Roughness, Metallic and Specular, ensuring compatibility with both Diffuse/Specular (e.g. [5]) and Base Color/Metallic (e.g. [28]) workflows. Some materials in our dataset also have an opacity map to handle, for example, surfaces with holes. Finally, we process the previously available dataset to enrich it with metallic, base color, and height properties, facilitating their combinations for a total of $\sim 6,000$ materials available for research, with clear licenses attached.
To validate the benefit of these additional data, we train recent methods with publicly available implementation and evaluate both qualitatively and quantitatively the results quality for different single image material capture methods [5, 43] and material generation [45], showing that the additional data enables higher quality acquisition and improved diversity in the generation results.
@@ 2.1 Existing dataset
The only large-scale, publicly available, spatially-varying material dataset contains 1628 distinct, pre-augmentation materials [5] with a research-only license. 512x512 [6] and 2K [7] versions of the materials were later published. It is based on random variations from a set of 155 procedural materials gathered from Allegorithmic Substance Share in 2017, distributed in 9 categories: paint(6), plastic(5), leather(13), metal(35), wood(23), fabric(6), stone(25), ceramic tiles(29), ground(13). Other datasets used in the field contain 100 and 62 measured BRDFs [29, 10], 170 SVBRDFs licensed from VRay[23, 46] or more recently 211 materials from online libraries [17]. Adobe 3D Assets [1]’s $11,000+$ procedural assets have been used to unlock significant quality improvement in the methods that could benefit from it for acquisition, generation, retrieval and captioning [50, 28, 44, 9]. This data is however difficult to license for Machine Learning, limiting its use to company internal users. In contrast, we propose a systematic gathering of non-duplicate, tileable, publicly available materials with permissive licenses from various online libraries, assembling $4,000$+ materials and their metadata, significantly increasing the variety and amount of ressources available for material-related research.
@@ 3 The Dataset
The MatSynth dataset is designed to support modern learning-based techniques for a variety of material-related tasks including, but not limited to, material acquisition, material generation, and synthetic data generation, e.g., for retrieval or segmentation. It improves on existing available datasets by providing a large collection ($\sim 2.5$ times larger than previously available datasets) of non-duplicate, high-quality, high-resolution realistic materials alongside useful metadata. Each material in the dataset is represented by a set of reflectance maps (basecolor, diffuse, normal, height, roughness, metallic, and specular), complemented by fine-grained annotations including the source and link to the asset, its license, and original author (when available), tags, the creation method, whether is stationary or not, a timestamp for versioning, a description (when available), and the material physical size (when available).
@@ 3.3 Data Processing
Finally, we automatically compute an estimate of the correct displacement factor for each material, by optimizing a displacement factor parameter to match the normal vectors derived from the scaled height map to the ground truth normal map. This ensures a good match between the normal and the displacement scale when rendering. We provide the correctly scaled height map, as well as the original and the scaling parameters (scale factor and mid-value compatible with the Blender “Displacement” node).
@@ 3.4 Data Augmentation
Rotation and crops and illumination. The collected 4069 materials are augmented by rotation and cropping. In particular, we extract 168 unique crops for each material, at different scales. The crops are extracted in the following way. The original material is rotated 8 times ($0\degree$, $45\degree$, $90\degree$, $135\degree$, $180\degree$, $225\degree$, $270\degree$, $315\degree$)–for angles that are not multiple of $90\degree$, we first tile the material, apply the rotation, and finally crop the center of the material at 4K resolution. For each rotation, we extract 16 non-overlapping crops at 1K resolution, 4 at 2K resolution, 1 at 4K resolution, for a total of 21 crops per rotation. All crops are resized to a resolution of $1024\times 1024$ pixels. Each crop is rendered under different environment illuminations (2 outdoors, 2 indoors, and 1 studio light), for a total of 840 renders per material. For each render, a random rotation is applied to the environment map. The resulting dataset consists of 683,592 material crops and 3,417,960 renders. Fig. 2 shows the five renderings under different environment illumination for 4 different materials from the dataset. We include the 5 environment maps used for rendering in the supplemental materials.
Rendering. We render each material displacing the mesh using the available Height map, thus introducing realistic cast shadows in the rendering. However, to maintain a pixel match with the maps, required for supervised SVBRDF estimation tasks, we opt for an orthographic rendering. This introduces yet another issue, as specular reflections are lost on perfectly flat surfaces due to rays being parallel when rendering with an orthographic camera. To restore specular reflection while keeping pixel match and cast shadows, we split the rendering in two separate passes. Similar to the data used to train ControlMat [44], we first render the diffuse component of the material using an orthographic camera with displacement enabled, we then render the specular component using a perspective camera, while disabling the displacement. The two renders are then added back to get the final render. Fig. 3 shows the output of the two render passes and the final result. We evaluate each augmentation in an ablation study in the Supplemental Materials.
@@
Figure 3: Render samples using the two-pass strategy. This ensures that the maps and the rendering are well aligned, avoiding parrallax effects, but preserving specular highlights.
@@ 3.6 Dataset Statistics
We report here statistics of our dataset in terms of categories, tags, creation methodology, stationarity, and number of asset per source.
Categories. We report the category histogram in Fig. 4, showing a wide diversity of appearances, and compensating for some gaps in the distribution of previously available materials.
Tags. We have $21,737$ tags from $1,239$ unique tags associated with the materials. For example, the top five most represented tags are floor(924), dirty(543), man (504), wood (474), brown (444) and wall(424). Each material has between 0 (371 assets didn’t have associated tags) and 20 tags with an average of 5.74, a standard variation of 4.35 and a median of 5.
@@
Stationarity. Our dataset contains 3061 stationary and 1008 non-stationary materials. We compute stationarity by extracting 8 random crop of the basecolor and height map and comparing the CLIP cross-similarity matrix. If more than 12.5% of the cross-similarities are below a threshold (we empirically found 0.9 to work well), the material is not considered stationary.
@@
[…]
@@
[…]
@@ 4.1 Test dataset
We build a static test set, with the aim of simplifying the comparison of future work, by handpicking 5 unique samples for each category. We further ensure that none of the chosen samples is a simple variation (e.g.: color or scale) of other items in the training set to avoid any bias. We merge this dataset, consisting of 65 materials, with the test set from Deschaintre et al. [7], for a total of 89 test materials.
@@ 4.2 Evaluation Methodology
Evaluation is carried on two main materials–related tasks: estimation and generation. For both tasks, we look at methods with publicly available implementations. In particular, we evaluate Deschaintre et al. [5] and SurfaceNet [43] for materials capture performance, and MatFuse [45] for material generation. In both cases, we focus on the diffuse/specular rendering workflow for an easier comparison with the current state of the art. We additionally propose an evaluation of SurfaceNet trained using the basecolor/metallic workflow in Supplemental material, to set a new baseline for future work shifting to this workflow.
@@ 4.3 Training Strategy
All methods are trained on PC with a single NVIDIA RTX 4090 GPU. Training of Deschaintre et al. [5] and SurfaceNet [43] is carried out on a flash-lit version of the dataset, as these methods are designed for SVBRDF estimation from flash-lit pictures.
Deschaintre et al. (2018) We train the method, using the author’s publicly available Tensorflow implementation, on a combination of our dataset and the original one. Following the original paper, we train the model for $400,000$ steps (72 hours) with a batch size of 8 and a learning rate of $2^{-5}$.
SurfaceNet Similarly to the training of Deschaintre et al. [5], we use the author’s original Pytorch implementation and the combined dataset. We also follow the original paper and train the model for $400,000$ steps (96 hours) with a batch size of 8 and a learning rate of $2^{-5}$. We include the adversarial discriminator after $100,000$ steps.
Diffusion Model We evaluate the quality improvement of generation provided by our dataset on the recent MatFuse [45]. We evaluate the performance both when fine-tuning the diffusion model and when training from scratch. For the first experiment, we use the same multi-encoders compression model, pre-trained on the Deschaintre et al. [5]’s dataset, but fine-tune the latent diffusion model itself. In particular, the diffusion model is fine-tuned for 150,000 iterations with a batch size of $20$ and a learning rate of $10^{-5}$. We use a linear learning rate warm-up until step 10,000.
For the second experiment, we train the entire model, on our dataset. In particular, we train the autoencoder model for 1.000.000 iterations with a batch size of $4$ and a learning rate of $10^{-4}$. We include the adversarial discriminator after $300,000$ steps. Finally, we train the diffusion model for $500,000$ iterations with a batch size of $16$ using an AdamW [27] optimizer, with a learning rate value of $10^{-4}$, with a linear learning rate warm-up until step 10,000.
For inference, following the original paper, we denoise using a DDIM sampling schedule [39] with $50$ timesteps.
@@ 4.4 SVBRDF Estimation Results
Figure 7: Comparison on Material Generation. We show renderings of 10 randomly generated materials, both with and without our dataset. These results are directly sampled from the network and have not been filtered. As expected, additional data help achieve better and more diverse results.
@@ 4.5 Materials Generation Results
We evaluate material generation using MatFuse [45], trained on Deschaintre et al. [5]’s dataset only, and comparing it against a version fine-tuned with MatSynth and one fully trained on it. As material maps have different distributions than natural images, we report FID on renderings under environment lighting. For fairness, we only use Deschaintre et al. [5]’s dataset to compute the Ground-Truth distribution and compute the FID for each method using $1,000$ random samples. Before fine-tuning we find a FID score (lower is better) of $239.9$, $210.3$ after fine-tuning, and $89.84$ when training from scratch. We believe the performance difference to be caused by the retraining of the autoencoder using MatSynth, while the other experiments used the original VQ-VAE from MatFuse, trained on Deschaintre et al. [5]’s dataset only.
We further evaluate the benefit of the additional data for diversity and quality. We show in Table 3 the percentage of samples belonging to different classes, and show a qualitative comparison in Fig. 7. Both evaluations show a wider diversity of results, and Fig. 7 shows that the generation results are more realistic after fine-tuning or when training on MatSynth from scratch.
original fine-tuned full training
Wood 36.1% 25.72% 7.5%
Ceramic 1.1% 12.51% 8.5%
Stone 21.5% 14.8% 7.3%
Metal 3.45% 8.4% 6.4%
Fabric 0.33% 7.34% 7.8%
Ground 1.8% 16.2% 6.4%
Others 35.72% 15.03% 56.1%